OpenStudy (anonymous):

PLEASE HELP! WILL GIVE MEDAL! For the function 3x – 2y = 12, what is the x-intercept? -4 4 6 -6

OpenStudy (anonymous):

first solve for x

OpenStudy (anonymous):

2? @maths111

OpenStudy (anonymous):

Translate the expression into a word phrase. k + 12 A. a number less than 12 B. 12 more than a number C. the product of a number and 12 D. the quotient of a number and 12

OpenStudy (anonymous):

Do you know ,how to solve for x?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

12 more than a number

OpenStudy (anonymous):

3x -2y =12 3x = 12 + 12

OpenStudy (anonymous):

3x = 12 +2y

OpenStudy (anonymous):

3x/3 = 12/3 + 2y/3

OpenStudy (anonymous):

So what is the answer?

OpenStudy (anonymous):

x = 4 + 2y/3 so you answer is 4 . x = my +c

OpenStudy (anonymous):

Im confused, sorry.

OpenStudy (anonymous):

wait....

OpenStudy (anonymous):